Circular No. 6822

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1630–47 E. Kuulkers, Oxford University; S. Dieters, University of Alabama,

Huntsville; M. McCollough, Universities Space Research Association and Marshall Space Flight Center; M. Wieringa, Australia Telescope National Facility, Narrabri; M. Buxton, Mount Stromlo and Siding Spring Observatories, Australian National University; and R. Remillard, Massachusetts Institute of Technology (MIT), on behalf of the RXTE/ASM team at MIT and Goddard Space Flight Center, NASA, report: “The RXTE ASM (2–12 keV) shows that the recurrent blackhole x-ray transient 4U 1630–47 (cf. IAUC 6389) has started to become active again, as predicted by Kuulkers et al. (1997, MNRAS 291, 81). It rose to ∼ 40 mCrab between Feb. 1.0 and 2.7 UT, where it stayed more or less constant for the next ∼ 6 days. It then rose to ∼ 60 mCrab on Feb. 8.3 and ∼ 90 mCrab on Feb. 9.3. Subsequent RXTE PCA observations on Feb. 9.7 show the source at 140 mCrab (2–60 keV). More RXTE observations have been scheduled. GRO BATSE has also detected 4U 1630–47 between Feb. 1 and 6 with a maximum flux of 360 mCrab (20–100 keV) with a photon index near −2.5. ATCA observations at 8640 MHz on Feb. 9.1 and 9.8 failed to detect a radio counterpart and set an upper limit of 0.7 mJy. This source has shown similar x-ray properties to the Galactic superluminal sources GRO J1655–40 and GRS 1915+105 and so may show relativistic ejections during its outburst. No radio or optical counterpart has been identified yet for this source, and we therefore urge observations at other wavelengths. More information on 4U 1630–47 can be found at the URL http://www-

astro.physics.ox.ac.uk/~erik/1630/index.html.”

1SAX J0054.9–7226 G. L. Israel, Osservatorio Astronomico di Roma, communicates on be-

half of the BeppoSAX Science Team: “The position of the XTE/BeppoSAX 59-s transient x-ray pulsar in the Small Magellanic Cloud, reported on IAUC 6818, is also included in an archival ROSAT HRI observation obtained on 1996 Apr. 26. The source is detected at α = 0^{h}54^{m}56^{s}.1, δ = −72^{o}26′45′′ (equinox 2000; uncertainty radius 10′′) with a 0.1–2.4-keV flux of ∼ 1.6 × 10^{−13} erg s^{−1} cm^{−2}. Visual inspection of European Southern Observatory plates reveals three unclassified stars (V > 15.5) within the refined error circle. Optical follow-up observations of these candidates are strongly encouraged.”

1998 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green